An satronaut applies a net force of 8000 N to an asteriod and it accelerates at 14.7 m/s^2. What is the asteriod's mass to two sig figs?
Answer #1

Using Newton's 2nd law

F = ma

m = F/a = 8000N/ 14.7 m/s^2= 544.22 kg ans

rate me if it was helpful. Thanks in advance